# Beyond the Leaderboard: Rethinking Medical Benchmarks for Large Language Models

### v1

@@ Abstract
Large language models (LLMs) show significant potential in healthcare, prompting numerous benchmarks to evaluate their capabilities. However, concerns persist regarding the reliability of these benchmarks, which often lack clinical fidelity, robust data management, and safety-oriented evaluation metrics. To address these shortcomings, we introduce MedCheck, the first lifecycle-oriented assessment framework specifically designed for medical benchmarks. Our framework deconstructs a benchmark’s development into five continuous stages, from design to governance, and provides a comprehensive checklist of 46 medically-tailored criteria. Using MedCheck, we conducted an in-depth empirical evaluation of 53 medical LLM benchmarks. Our analysis uncovers widespread, systemic issues, including a profound disconnect from clinical practice, a crisis of data integrity due to unmitigated contamination risks, and a systematic neglect of safety-critical evaluation dimensions like model robustness and uncertainty awareness. Based on these findings, MedCheck serves as both a diagnostic tool for existing benchmarks and an actionable guideline to foster a more standardized, reliable, and transparent approach to evaluating AI in healthcare.
@@
Zizhan Ma1  Wenxuan Wang2  Guo Yu3  Yiu-Fai Cheung 1 Meidan Ding3  Jie Liu4  Wenting Chen4  Linlin Shen3 1 The Chinese University of Hong Kong   2 Renmin University of China 3 Shenzhen University   4 City University of Hong Kong 1zzma2@cse.cuhk.edu.hk2wangwenxuan@ruc.edu.cn3wentichen7-c@my.cityu.edu.hk
@@ 1 Introduction
However, despite their widespread adoption, the reliability of these benchmarks is a growing concern. Many evaluations rely heavily on closed-form, multiple-choice questions (MCQA) that, while testing factual knowledge, fail to assess open-ended clinical reasoning or account for real-world variability (Wu et al., 2025a; Zhang et al., 2025). Furthermore, a significant number of benchmarks are constructed from academic materials rather than authentic clinical data, which raises critical concerns about their clinical fidelity and the risk of data contamination—where models are evaluated on data seen during training (Ouyang et al., 2024; Wu et al., 2025b). These systemic flaws can create an illusion of progress, where models are optimized for tasks that lack genuine clinical utility.
Recent efforts have sought to improve evaluation quality by proposing general frameworks for AI benchmarks, such as BetterBench for general-purpose AI (Reuel et al., 2024) and How2Bench for code (Cao et al., 2025). While these frameworks have uncovered systemic issues in data quality and reproducibility, they are not tailored to the unique, high-stakes requirements of medicine. The medical domain demands a more rigorous, context-aware approach that accounts for specialized terminology, the ethical handling of sensitive patient data, and the paramount importance of safety and reliability.
To develop trustworthy clinical AI, the field must therefore transition from an ad-hoc, publication-driven approach to a disciplined, engineering-oriented paradigm for evaluation (Laskar et al., 2024; Yan et al., 2024). Mature fields like software engineering have long relied on structured methodologies such as the Software Development Life Cycle (SDLC) to manage complexity and ensure quality (Yas et al., 2023; Royce, 1970). Adopting a similar lifecycle-aware perspective is not merely beneficial for the science of AI evaluation; it is an absolute prerequisite for ensuring safety and efficacy in the medical domain (Park et al., 2020).
In this paper, we introduce MedCheck, the first comprehensive, lifecycle-oriented assessment framework specifically engineered for medical LLM benchmarks. MedCheck deconstructs benchmark development into five continuous stages, from design to governance, and provides a checklist of 46 medically-tailored criteria. To demonstrate its utility and diagnose the state of the field, we applied MedCheck in an in-depth empirical evaluation of 53 existing medical benchmarks. Our analysis reveals widespread, systemic issues: a profound disconnect from clinical practice, a crisis of data integrity from unmitigated contamination risks, and a systematic neglect of safety-critical dimensions like model robustness and uncertainty awareness.
Our contributions are threefold:
@@
  • We conduct an in-depth evaluation of 53 medical benchmarks, revealing widespread, systemic weaknesses across the current evaluation landscape.
@@ 2.1 The Evolution of Medical LLM Benchmarks
To bridge this gap, recent efforts have focused on creating benchmarks that better mirror real-world clinical tasks. These include comprehensive benchmarks covering report summarization, diagnosis, and treatment planning (Wu et al., 2025b; Liu et al., 2024a; Wu et al., 2025a), as well as agentic benchmarks that simulate sequential clinical decision-making (Jiang et al., 2025; Schmidgall et al., 2024). Although these newer benchmarks represent a significant step forward, their development has often outpaced the creation of rigorous standards for their own evaluation, a gap our work aims to fill.
@@ 2.2 Towards a Science of Benchmark Evaluation
The proliferation of benchmarks has ignited a critical discussion about their reliability and validity. Alaa et al. (2025) empirically demonstrated a poor correlation between LLM performance on existing benchmarks and in real-world clinical scenarios, highlighting a crisis in the construct validity of current evaluation methods—the degree to which a test measures what it claims to measure (Cronbach and Meehl, 1955). This has catalyzed a new research area focused on the science of benchmarking itself.
@@
Pioneering works have proposed general frameworks to standardize the evaluation of AI benchmarks. For instance, Bhardwaj et al. (2024) developed a checklist for data curation best practices. More comprehensively, BetterBench (Reuel et al., 2024) provides a 46-criteria framework for general-purpose AI benchmarks, and How2Bench (Cao et al., 2025) offers a 55-item checklist for code-related LLM benchmarks. These frameworks promote a lifecycle-aware perspective, revealing common issues in data quality, reproducibility, and transparency.
However, these valuable solutions are not specialized for medicine. High-stakes clinical applications demand a more rigorous, context-aware approach that accounts for specialized terminology, patient data ethics, and the paramount importance of safety and reliability. While existing surveys have profiled the medical LLM landscape (Liu et al., 2024c; Yan et al., 2024), they stop short of providing a structured framework for assessing benchmark quality. Our work bridges this critical gap by introducing MedCheck, the first assessment framework specifically engineered for the unique complexities of the medical domain.
@@ 3.1 The Five-Phase Medical Benchmark Lifecycle
This foundational phase moves beyond mere task definition to establish the benchmark’s theoretical underpinnings, focusing on its construct validity—the degree to which it accurately measures its intended theoretical construct, such as "clinical reasoning" (Cronbach and Meehl, 1955; Mehandru et al., 2025; Schmidgall et al., 2024). As forcefully argued by Alaa et al. (2025), high performance on exam questions may not correlate with real-world clinical utility. This phase, therefore, defines the benchmark’s specific, clinically relevant purpose and its intended contribution, setting the standard against which its validity will later be judged.
Figure 3: The three-step methodology employed in this study. Our approach involved (1) developing the 46-criteria assessment framework, (2) systematically curating and scoring 53 benchmarks against it, and (3) performing a quantitative synthesis to generate systemic insights.
@@
Phase V: Documentation, Openness, and Governance. This community-facing phase ensures the benchmark’s long-term value and impact. It involves clear documentation and open-source principles for transparency and reproducibility (Arvan et al., 2022; Cohen et al., 2018). Critically, it also demands a robust governance model for long-term maintenance, versioning, and community feedback to prevent the benchmark from becoming a static, decaying artifact.
@@ 3.2 Study Design
Our study employed a meticulous, multi-step methodology, illustrated in Figure 3, to ensure objectivity, reproducibility, and depth.
Step 1: Framework Development. We developed the 46-criteria MedCheck framework, which serves as the analytical lens for this study. The creation of this framework was a multi-stage process. First, we conducted an extensive literature review of existing best practices in general AI benchmarking (Reuel et al., 2024; Cao et al., 2025) and machine learning data curation (Bhardwaj et al., 2024). Second, we distilled these principles into a draft checklist. Finally, this checklist was iteratively refined into the final 46-criteria framework, grounded in the unique ethical and practical constraints of the medical domain, such as patient privacy (e.g., HIPAA compliance), the need for evidence-based standards, and the high stakes associated with potential patient harm (Farhud and Zokaei, 2021; Eke and Shuib, 2025). The complete 46-criteria checklist is provided in Appendix C.
Step 2: Systematic Benchmark Curation and Analysis. We curated and analyzed a corpus of 53 medical LLM benchmarks, listed in Appendix A. The selection process involved a systematic search for papers published between 2018-2025 using keywords such as "medical LLM," "medical benchmark," and "healthcare," aiming for comprehensive coverage of tasks, languages, and publication years. To ensure objectivity, a rigorous scoring protocol was implemented. Each benchmark was independently scored by a team of three NLP researchers with clinical informatics experience against the 46 criteria on a 3-point Likert scale (0: not met; 1: partially met; 2: fully met). All assessments were based exclusively on publicly available artifacts, including published papers, code repositories, and official websites. Any scoring discrepancies were resolved through a consensus discussion to arrive at a final score for each criterion.
Step 3: Quantitative Synthesis and Insight Generation. The individual scores were aggregated to enable a multi-level quantitative analysis. We calculated per-criterion average scores to identify specific, widespread weaknesses across the field. These were then rolled up into per-phase scores to assess the maturity of each stage in the development lifecycle. Finally, an overall score was computed for each benchmark to gauge its overall quality. This quantitative synthesis allowed us to move beyond anecdotal critiques and identify systemic, statistically significant deficiencies across the entire landscape of medical LLM evaluation, forming the evidence base for the guidelines presented in the following section. A detailed breakdown of these quantitative results can be found in Appendix B.
@@ 4 MedCheck: A Guideline for Medical Benchmarks
Based on our five-phase lifecycle and extensive analysis, we present the MedCheck framework as an actionable guideline. For each phase, we explain its core principle and summarize our key findings, further detailed in Appendix B.
@@ 5 Does it define the scope of the medical specialties of the benchmark?
6 Is it designed to meet the needs of LLM researchers and clinical challenges?
@@ 4.1 Guideline for Design & Conceptualization
Explanation. This foundational phase anchors a benchmark in clinical relevance and scientific novelty. It necessitates the precise definition of the targeted LLM capabilities, medical scope, and application value, all of which must be grounded in documented domain expertise and authoritative sources. A robust design requires alignment with established medical standards (e.g., ICD), the use of clearly defined, multi-dimensional metrics that extend beyond accuracy, and the proactive consideration of safety and bias from inception (Arora et al., 2025).
Findings: The Clinical Disconnect. Our analysis reveals that while nearly all benchmarks (98%) define high-level objectives, this superficial clarity masks a systemic issue we term the Clinical Disconnect. A concerning 53% (28 of 53) fail to align with any formal medical standards (e.g., ICD, SNOMED CT), isolating them from the structured language of clinical practice. Furthermore, 47% (25 of 53) do not incorporate safety and fairness into their design, and 36% (19 of 53) evaluate only a single dimension like accuracy, neglecting critical aspects such as completeness. This disconnect stems from an "academic-first, clinical-second" mindset, where developers favor convenient data sources like exam questions from MedQA (Jin et al., 2021) and MedMCQA (Pal et al., 2022) over data reflecting complex clinical workflows. This convenience-driven design compromises clinical fidelity and construct validity (Alaa et al., 2025), leading to models optimized for irrelevant tasks—a perilous path that produces tools unfit and potentially unsafe for real-world deployment.
@@ 4.2 Guideline for Dataset Construction & Managemen
Findings: A Crisis of Foundational Validity. Our analysis reveals critical weaknesses in data management that undermine the field’s empirical foundation. While most benchmarks are reasonably transparent about their primary data sources, subsequent quality control is severely lacking. A staggering 92% (49 of 53) fail to address data contamination, often by directly reusing data from other benchmarks without verification. Furthermore, 70% (37 of 53) are insufficiently diverse or representative, and 58% (31 of 53) lack any clear data audit or expert review mechanism. This widespread failure to ensure data integrity poses a direct threat to the field’s validity. Unvetted data may contain factual errors, while contamination leads to artificially inflated scores, creating a false perception of model capability (Magar and Schwartz, 2022; Deng et al., 2024; Bender et al., 2021). This renders leaderboards misleading and undermines the scientific credibility of reported results (Holistic AI, 2024), risking the construction of a field built upon unreliable evidence.
@@ 4.3 Guideline for Technical Implementation & Evaluation Methodology
Findings: Systematic Neglect of Safety-Critical Capabilities. This phase is the most underdeveloped in our analysis (average score: 52.4%), revealing a profound gap between current practice and the requirements for reliable medical AI. An alarming 94% (50 of 53) of benchmarks have no mechanism to test for model robustness, and 96% (51 of 53) fail to evaluate a model’s ability to handle uncertainty. Furthermore, 51% (27 of 53) neglect the model’s reasoning process, focusing only on the final answer. These omissions constitute a systematic neglect of safety. A model’s reasoning, robustness, and uncertainty awareness are cornerstones of clinical trustworthiness (Farhud and Zokaei, 2021). A brittle model that fails with slight data variations is dangerous (The BMJ, 2024), and an overconfident model that gives incorrect advice without expressing uncertainty is a direct threat to patient safety. By not measuring these capabilities, the community implicitly deems them unimportant, increasing the risk of deploying brittle, opaque, and unsafe systems.

### v2

@@ Abstract
Large language models (LLMs) show significant potential in healthcare, prompting numerous benchmarks to evaluate their capabilities. However, concerns persist regarding the reliability of these benchmarks, which often lack clinical fidelity, robust data management, and safety-oriented evaluation metrics. To address these shortcomings, we introduce MedCheck, the first lifecycle-oriented assessment framework designed for medical benchmarks. Our framework deconstructs benchmark development into five stages from design to governance, and provides a comprehensive checklist of 46 medically-tailored criteria. Using MedCheck, we conducted an in-depth empirical evaluation of 56 medical LLM benchmarks. Our analysis uncovers widespread, systemic issues, including a profound disconnect from clinical practice, a crisis of data integrity due to unmitigated contamination risks, and a systematic neglect of safety-critical evaluation dimensions like model robustness and uncertainty awareness. Based on these findings, MedCheck serves as a diagnostic framework to audit existing benchmarks and an actionable guideline for a more standardized, reliable, and transparent approach to evaluating AI in healthcare.
@@
Wenting Chen1   Guo Yu2   Yiu-Fai Cheung3   Meidan Ding2 Jie Liu4Zizhan Ma3thanks:   Zizhan Ma is the corresponding author.Wenxuan Wang5Linlin Shen2 1 Stanford University  2 Shenzhen University  3 The Chinese University of Hong Kong 4 City University of Hong Kong  5 Renmin University of China wentchen@stanford.eduwangwenxuan@ruc.edu.cnzizhan.ma@link.cuhk.edu.hk
@@ 1 Introduction
However, despite their widespread adoption, the reliability of many newer benchmarks is a growing concern. Echoing critiques from within the clinical informatics community (Wornow et al., 2023), many evaluations rely heavily on closed-form, multiple-choice questions (MCQA) that, while testing factual knowledge, fail to assess open-ended clinical reasoning or account for real-world variability (Wu et al., 2025a; Zhang et al., 2025). Furthermore, a significant number are constructed from academic materials rather than authentic clinical data, which raises critical concerns about their clinical fidelity and the risk of data contamination—where models are evaluated on data seen during training (Ouyang et al., 2024; Wu et al., 2025b). These systemic flaws, particularly prevalent in rapidly developed benchmarks designed for general-purpose LLMs, can create an illusion of progress where models are optimized for tasks that lack genuine clinical utility, a problem long recognized by researchers working with real-world electronic health record (EHR) data (Wornow et al., 2023; Blagec et al., 2023; Alaa et al., 2025).
Recent efforts have proposed general frameworks to improve evaluation quality, such as BetterBench for general-purpose AI (Reuel et al., 2024) and How2Bench for code (Cao et al., 2025). However, while valuable, these frameworks are not tailored for the medical domain, which demands a more rigorous approach accounting for specialized terminology, patient data ethics, and the paramount importance of safety.
To develop trustworthy clinical AI, the field must transition from an ad-hoc, publication-driven approach to a disciplined, engineering-oriented paradigm for evaluation (Laskar et al., 2024; Yan et al., 2024). Adopting a lifecycle-aware perspective, similar to principles in mature engineering fields, is a prerequisite for ensuring safety and efficacy in the medical domain (Park et al., 2020).
We introduce MedCheck, the first comprehensive, lifecycle-oriented assessment framework for medical LLM benchmarks. MedCheck deconstructs benchmark development into five continuous stages, from design to governance, and provides a checklist of 46 medically-tailored criteria. We demonstrate MedCheck’s utility by applying it in an in-depth evaluation of 56 medical benchmarks. Our analysis reveals widespread, systemic issues: a profound disconnect from clinical practice, a crisis of data integrity from unmitigated contamination risks, and a systematic neglect of safety-critical dimensions like model robustness and uncertainty awareness. Our contributions are threefold:
@@
  • We conduct an in-depth evaluation of 56 medical benchmarks, revealing widespread, systemic weaknesses across the current evaluation landscape.
@@ 2.1 The Evolution of Medical LLM Benchmarks
To bridge this gap, recent efforts include data-centric frameworks like BigBIO (Fries et al., 2022) and public clinical datasets like MIMIC-IV (Johnson et al., 2023). There is also a growing emphasis on creating benchmarks that better mirror real-world clinical tasks, such as comprehensive benchmarks that cover report summarization, diagnosis, and treatment planning (Wu et al., 2025b; Liu et al., 2024a; Wu et al., 2025a; Bedi et al., 2025), as well as agentic benchmarks that simulate sequential clinical decision-making (Jiang et al., 2025; Schmidgall et al., 2024). Although these efforts represent a significant step forward, their development has often outpaced the creation of rigorous standards for their own evaluation, a gap our work aims to fill.
@@ 2.2 Towards a Science of Benchmark Evaluation
The proliferation of benchmarks has ignited discussion about their reliability and validity. Alaa et al. (2025) empirically demonstrated a poor correlation between LLM performance on existing benchmarks and in real-world clinical scenarios, highlighting a crisis in the construct validity of benchmarks—the degree to which a test measures what it claims to measure (Cronbach and Meehl, 1955). Wornow et al. (2023) also finds that foundation modals are often evaluated on tasks that do not provide meaningful insights on their usefulness to health systems. This has catalyzed a new research area focused on the science of benchmarking itself.
@@
Pioneering works have proposed general frameworks to standardize the evaluation of AI benchmarks. For instance, Bhardwaj et al. (2024) developed a checklist for data curation best practices. More comprehensively, BetterBench (Reuel et al., 2024) provides a 46-criteria framework for general-purpose AI benchmarks, and How2Bench (Cao et al., 2025) offers a 55-item checklist for code-related LLM benchmarks. These frameworks promote a lifecycle-aware perspective, revealing common issues in data quality, reproducibility, and transparency. There are also initiatives like TRIPOD-LLM (Gallifant et al., 2025) which propose standardized reporting guidelines to enhance transparency and reproducibility.
However, these valuable solutions are either general-purpose or focused on reporting. High-stakes clinical applications demand a more rigorous, context-aware approach that accounts for specialized terminology, patient data ethics, and the paramount importance of safety and reliability. While existing surveys have profiled the medical LLM landscape (Liu et al., 2024c; Yan et al., 2024), they stop short of providing a structured framework for assessing benchmark quality. Our work bridges this critical gap by introducing MedCheck, the first assessment framework specifically engineered for the unique complexities of the medical domain.
For a detailed discussion of the foundational evaluation and reporting frameworks of the clinical informatics community, refer to the Appendix A.
@@ 3.1 The Five-Phase Medical Benchmark Lifecycle
This foundational phase moves beyond mere task definition to establish the benchmark’s theoretical underpinnings, focusing on its construct validity—the degree to which it accurately measures its intended theoretical construct, such as "clinical reasoning" (Cronbach and Meehl, 1955; Mehandru et al., 2025; Schmidgall et al., 2024). This phase defines the benchmark’s specific, medically relevant purpose and intended contribution, as exam performance may not correlate with real-world utility (Alaa et al., 2025).
Figure 3: The three-step methodology employed in this study. Our approach involved (1) developing the 46-criteria assessment framework, (2) systematically curating and scoring 56 benchmarks against it, and (3) performing a quantitative synthesis to generate systemic insights.
@@
Phase V: Documentation, Openness, and Governance. This community-facing phase ensures the benchmark’s long-term value and impact. It involves clear documentation and open-source principles for transparency and reproducibility (Arvan et al., 2022; Cohen et al., 2018). Critically, it demands a robust governance model for long-term maintenance, versioning, and community feedback.
MedCheck evaluates benchmarks as complete engineered systems, not just data collections. It recognizes that even when datasets are reused, significant value can be added through innovative task design, rigorous validation, and sustainable governance.
@@ 3.2 Study Design
Our three-step methodology (Figure 3) was designed to ensure objectivity, reproducibility, and depth.
Step 1: Framework Development. We developed the 46-criteria MedCheck framework, which serves as the analytical lens for this study. We developed the framework by first conducting an extensive literature review of existing best practices in general AI benchmarking (Reuel et al., 2024; Cao et al., 2025) and machine learning data curation (Bhardwaj et al., 2024), which we then distilled and iteratively refined into the final 46-criteria framework, grounded in the medical domain’s unique ethical and practical constraints, such as patient privacy (e.g., HIPAA compliance), the need for evidence-based standards, and the high stakes associated with potential patient harm (Farhud and Zokaei, 2021; Eke and Shuib, 2025). The complete 46-criteria checklist is provided in Appendix H.
Step 2: Systematic Benchmark Curation and Analysis. We curated and analyzed a corpus of 56 medical LLM benchmarks, listed in Appendix E. The selection process is described in Appendix B To ensure objectivity, a rigorous scoring protocol was implemented. Our evaluation process incorporates both LLM-as-judge and domain expert review to balance scalability with accuracy. Specifically, we first provided the paper text and documentation of each benchmark to a large language model and requested an initial scoring based on our 46 criteria. Then, a panel of three NLP researchers with clinical informatics experience reviewed and adjusted these scores on a 3-point Likert scale (0: not met; 1: partially met; 2: fully met). All assessments were based exclusively on publicly available artifacts, including published papers, code repositories, and official websites. Any scoring discrepancies were resolved through a consensus discussion to arrive at a final score for each criterion. We provide the detailed scoring guidelines for the three NLP researcher annotators, along with the full human-in-the-loop annotation protocol, in Appendix C.
Step 3: Quantitative Synthesis and Insight Generation. The individual scores were aggregated to enable a multi-level quantitative analysis. We calculated per-criterion average scores to identify specific, widespread weaknesses across the field. These were then rolled up into per-phase scores to assess the maturity of each stage in the development lifecycle. Finally, an overall score was computed for each benchmark to gauge its overall quality. This quantitative synthesis allowed us to move beyond anecdotal critiques and identify widespread, descriptive patterns of deficiencies across the entire landscape of medical LLM evaluation, forming the evidence base for the guidelines presented in the following section. A detailed breakdown of these quantitative results can be found in Appendix F.
@@ 4 MedCheck: A Guideline for Medical Benchmarks
Based on our five-phase lifecycle and extensive analysis, we present the MedCheck framework as an actionable guideline. For each phase, we explain its core principle and summarize our key findings, further detailed in Appendix F. Our analysis also differentiates between benchmarks designed for foundational medical knowledge and those for clinical practice. A detailed domain-specific analysis can be found in Appendix D, which shows that our framework appropriately evaluates benchmarks according to their intended purpose. In the tables that follow, each criterion number is hyperlinked to its full definition and scoring rubric in Appendix H.
@@ 5 Does it define the scope of the medical specialties of the benchmark?
6 Is it designed to meet the needs of LLM researchers and medical challenges?
@@ 4.1 Guideline for Design & Conceptualization
Explanation. This foundational phase anchors a benchmark in clinical relevance and scientific novelty. It necessitates the precise definition of the targeted LLM capabilities, medical scope, and application value, all of which must be grounded in documented domain expertise and authoritative sources. A robust design benefits from considering or aligning with established medical standards or terminologies (e.g., ICD, SNOMED CT) where appropriate to reflect real-world practice and facilitate downstream integration with clinical systems. It should also incorporate the use of clearly defined, multi-dimensional metrics that extend beyond accuracy, and the proactive consideration of safety and bias from inception (Arora et al., 2025).
Findings: The Clinical Disconnect. While nearly all benchmarks (98%) define high-level objectives, our analysis reveals a systemic issue we term the Clinical Disconnect. A total of 50% (28 of 56) fail to align with any formal medical standards (e.g., ICD, SNOMED CT). Furthermore, 45% (25 of 56) do not incorporate safety and fairness into their design, and 34% (19 of 56) evaluate only a single dimension like accuracy, neglecting critical aspects such as completeness. This disconnect stems from an "academic-first, clinical-second" mindset, where developers favor convenient data sources like exam questions from MedQA (Jin et al., 2021) and MedMCQA (Pal et al., 2022) over data reflecting complex clinical workflows. This convenience-driven design compromises clinical fidelity and construct validity, leading to models optimized for irrelevant and potentially unsafe tasks (Alaa et al., 2025).
@@ 4.2 Guideline for Dataset Construction & Managemen
Findings: A Crisis of Foundational Validity. Our analysis reveals critical data management weaknesses that undermine the field’s empirical foundation. While most benchmarks are reasonably transparent about their primary data sources, subsequent quality control is severely lacking. A staggering 88% (49 of 56) fail to address data contamination. While post-hoc detection is challenging for closed-source models, the field lacks proactive mitigation strategies within developer control, such as the implementation of canary strings or temporal data cutoffs to explicitly signal exclusion from future pre-training crawls. Furthermore, 66% (37 of 56) are insufficiently diverse or representative, and 55% (31 of 56) lack any clear data audit or expert review mechanism. This failure to ensure data integrity threatens the field’s validity. Unvetted data may contain factual errors, while contamination leads to artificially inflated scores, creating a false perception of model capability (Magar and Schwartz, 2022; Deng et al., 2024; Bender et al., 2021). This renders leaderboards misleading and undermines scientific credibility by building on unreliable evidence (Holistic AI, 2024).
@@ 4.3 Guideline for Technical Implementation & Evaluation Methodology
Findings: Systematic Neglect of Safety-Critical Capabilities. This is the most underdeveloped phase in our analysis (average score: 52.4%), revealing a profound gap between current practice and the needs of reliable medical AI. An alarming 89% (50 of 56) of benchmarks have no mechanism to test for model robustness, and 91% (51 of 56) fail to evaluate a model’s ability to handle uncertainty. Furthermore, 48% (27 of 56) neglect the model’s reasoning process, focusing only on the final answer. These omissions constitute a systematic neglect of safety. A model’s reasoning, robustness, and uncertainty awareness are cornerstones of clinical trustworthiness (Farhud and Zokaei, 2021). A brittle model that fails with slight data variations is dangerous (The BMJ, 2024), and an overconfident model that gives incorrect advice without expressing uncertainty is a direct threat to patient safety. By not measuring these capabilities, the community implicitly deems them unimportant, increasing the risk of deploying brittle, opaque, and unsafe systems.